Session: May 1998 (session 2)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (C. Bortolotti at Thu Jun 4 14:12:57 METDST 1998) Success
No known problems.

Torun (K. Borkowski at Thu Jun 4 16:10:54 METDST 1998) Success
No known problems

Westerbork (Tony Foley at Fri Jun 5 09:02:32 METDST 1998) Minor failures
I made up a procedure NULL to do nothing, and we had problems with the decoder during some parity checks. Otherwise looks OK.

Onsala (A. Polatidis at Mon Jun 8 17:18:17 METDST 1998) Minor failures
We lost the first hour of the experiment due to problems in the communication between the Field System and the Tape recorder.

Noto (C.Trigilio at Mon Jun 8 17:36:15 METDST 1998) Success

Jodrell (Alastair Gunn at Tue Jun 9 02:59:08 METDST 1998) Minor failures
IF1 recording Lovell LCP, IF2 recording Mk 2 LCP. Had to restart FS twice (23:41 and 23:53) to get correct tsys procedure working. Tsys okay from 00:01. Data missed were scans between 23:53 and 00:01.

Cambridge (Alastair Gunn at Tue Jun 9 02:59:34 METDST 1998) Minor failures
Low signal in BBC4 LSB due to microwave link limitations. Receiver cryostat alarm at about 00:30 UT - receiver began to warm, expect low signal.

Calibration (Jean\-Francois Desmurs at Wed Jun 10 15:52:27 METDST 1998) Minor failures
The ANTAB formatted file for this experiment can be found on VLBEER in the ? directory under the name er006.antab. Rem: Robledo: No Tsys. Read the comments in the file to get further information.